# Quantumness certification via non-demolition measurements

### v1

@@ IV.2 Three-measurement scheme
The simplest and most used version of the Leggett-Garg inequalities (LGIs) [5] is based on a three measurement scheme [10, 6]. To discuss the relations between the QNDM protocol, LGIs, and the macrorealism violation, we extend the protocol discussed in Sec. III to the three-measurement scheme depicted in Fig. 1. For technical details, the reader can refer to Ref. [26].
@@
Let us thus consider a system that evolves under a unitary transformation in the time interval $0\leq t\leq\mathcal{T}$, and the system observable $\hat{A}$ is measured sequentially at times $t_{0}\leq t_{1}\leq t_{2}=\mathcal{T}$. As above, the measurements of $\hat{A}$ are performed by coupling the system with a detector that interacts through the operator $\hat{u}$. The system time evolution in between the measurements is governed by the unitary operators $\hat{U}_{k}\equiv\hat{U}(t_{k},t_{k-1})\otimes\mathbbm{1}$ with $k=1,2$. Hence, overall, the time evolution operator $\hat{U}_{tot}$ for the system-detector within the interval $t_{0},t_{2}]$ read as
@@
$P_{\rm QNDM}(k,j,m,i,l)=U_{2,kj}U_{1,ji}\,\rho^{0}_{il}U^{*}_{1,lm}U^{*}_{2,mk}={\rm Tr}\left[\hat{\Pi}_{k}\hat{U}_{2}\hat{\Pi}_{j}\hat{U}_{1}\hat{\Pi}_{i}\,\hat{\rho}^{0}\hat{\Pi}_{l}\hat{U}^{\dagger}_{1}\hat{\Pi}_{m}\hat{U}^{\dagger}_{2}\right].$ (23)
are the probability amplitudes for the system to arrive in the state $\mathinner{|{k}\rangle}$ following the superposition of paths $i\rightarrow j\rightarrow k$ and $l\rightarrow m\rightarrow k$ (see Fig. 1).
Again, we can separate the classical $\mathcal{P}_{cl}(\Delta)$ from the quantum $\mathcal{P}_{q}(\Delta)$ contributions of $\mathcal{P}_{\rm QNDM}(\Delta)$. The first ones are obtained when $m=j$ and $l=i$,
$\mathcal{P}_{cl}(\Delta)=\sum_{i,j,k}P_{\rm QNDM}(k,j,j,i,i)\,\delta\left[\Delta-\Delta_{k,j,j,i,i}\right],$ (24)
and they are given by positive probabilities since $P_{\rm QNDM}(k,j,j,i,i)=P^{(0)}_{i}P^{(1)}_{i\rightarrow j}P^{(2)}_{j\rightarrow k}$ where $P^{(0)}_{i}=\rho^{0}_{ii}$ and $\{P^{(1)}_{i\rightarrow j}=|U_{1,ji}|^{2},P^{(2)}_{j\rightarrow k}=|U_{2,kj}|^{2}\}$ are the probabilities that the system is initially in the state $\mathinner{|{i}\rangle}$ and then makes a transition from state $\mathinner{|{i}\rangle}$ to $\mathinner{|{j}\rangle}$ and from $\mathinner{|{j}\rangle}$ to $\mathinner{|{k}\rangle}$, respectively.
The quantum contributions $\mathcal{P}_{q}(\Delta)$ of $\mathcal{P}_{\rm QNDM}(\Delta)$, instead, are present when $m\neq j$ or $l\neq i$, and thus they are associated with the coherent superposition of paths or trajectories in the space identified by the sequence of measurement outcomes [19, 26]. They formally read as
@@
IV.2.1 Noninvasive measurability condition
It can be checked that $\mathcal{P}_{\rm QNDM}(\Delta)$ in Eq. (22) satisfies the NSIT condition. That is, the marginals of $\mathcal{P}_{\rm QNDM}(\Delta)$, over all the measurement outcomes but the ones at a single time, return the ‘unperturbed’ probability to measure the observable $\hat{A}$  [^1] This probability is computed in correspondence of the current state of the system where the measurement is performed. given by the Born’s rule, without any intermediate measurement of the sequence affects the result.
@@ IV.2.1 Noninvasive measurability condition
$P(a_{k})$ $=$ $\sum_{i,l,j,m}P_{ND}(k,j,m,i,l)=\sum_{i,l,j,m}{\rm Tr}\left[\hat{\Pi}_{k}\hat{U}_{2}\hat{\Pi}_{j}\hat{U}_{1}\hat{\Pi}_{i}\,\hat{\rho}^{0}\hat{\Pi}_{l}\hat{U}^{\dagger}_{1}\hat{\Pi}_{m}\hat{U}^{\dagger}_{2}\right]=$ (26)
$=$ ${\rm Tr}\left[\hat{\Pi}_{k}\hat{U}\hat{\rho}^{0}\hat{U}^{\dagger}\right]={\rm Tr}\left[\hat{\Pi}_{k}\,\hat{\rho}(t_{2})\right],$
where here $\hat{U}=\hat{U}_{2}\hat{U}_{1}$ and $\hat{\rho}(t_{2})=\hat{U}\hat{\rho}^{0}\hat{U}^{\dagger}$. Therefore, $P(a_{k})$ is the probability to record the measurement outcome $a_{k}$ at time $t_{2}$ as given by Born’s rule [6].
We also compute the single-time probabilities $P(a_{i})$, $P(a_{l})$ for the initial measurement at time $t_{0}=0$:
@@
$=$ $\frac{1}{2}\sum_{l,j,m,k}\Big\{{\rm Tr}\left[\hat{\Pi}_{k}\hat{U}_{2}\hat{\Pi}_{j}\hat{U}_{1}\hat{\Pi}_{i}\,\hat{\rho}^{0}\hat{\Pi}_{l}\hat{U}^{\dagger}_{1}\hat{\Pi}_{m}\hat{U}^{\dagger}_{2}\right]+{\rm Tr}\left[\hat{\Pi}_{k}\hat{U}_{2}\hat{\Pi}_{j}\hat{U}_{1}\hat{\Pi}_{l}\,\hat{\rho}^{0}\hat{\Pi}_{i}\hat{U}^{\dagger}_{1}\hat{\Pi}_{m}\hat{U}^{\dagger}_{2}\right]\Big\}=$
$=$ ${\rm Tr}\left[\hat{\Pi}_{i}\,\hat{\rho}^{0}\right].$
@@
both these probabilities satisfy Born’s rule, as well as $P(a_{k})$.
@@
Moreover, for the measurement at time $t_{1}$, there are two possible realizations over the indices $j$ and $m$, so that
$P(a_{j})=\frac{1}{2}\sum_{i,l,m,k}\Big\{P_{ND}(k,j,m,i,l)+P_{ND}(k,m,j,i,l)\Big\}={\rm Tr}\left[\hat{\Pi}_{j}\hat{U}_{1}\hat{\rho}^{0}\hat{U}^{\dagger}_{1}\right]={\rm Tr}\left[\hat{\Pi}_{j}\,\hat{\rho}(t_{1})\right]$ (29)
@@
that is to the probability of recording the outcome $a_{j}$ after the system has evolved to the state $\hat{\rho}(t_{1})=\hat{U}_{1}\,\hat{\rho}^{0}\hat{U}^{\dagger}_{1}$, without the initial measurement at time $t=0$ has affected the system state.
@@
Finally, if we sum the QNDM quasi-probabilities $P_{\rm QNDM}(k,j,m,i,l)$ over three indices return the Kirkwood-Dirac quasi-probabilities at two-times [48, 25, 63], apart from the 3-tuples $(j,m,k)$ and $(i,l,k)$ that give rise to a single-time probability. By definition, Kirkwood-Dirac quasi-probabilities respect the NSIT assumption.
IV.3 Necessary and sufficient condition for the violation of macreorealism
@@
It can be proved [26] that a system is in a coherent superposition of the measured observable’s eigenstates if and only if the QNDM quasi-probability density function $\mathcal{P}_{\rm QNDM}(\Delta)$ exhibits negativity, i.e., it has negative contributions. The fact that the measured system is in a superposition state violates the MRps and, consequently, the MR. Therefore, the statement above can also be formulated as: the MR is violated if and only if $\mathcal{P}_{\rm QNDM}(\Delta)$ has negative contributions.
@@
This result is based on the following three main properties of the QNDM protocol. First, the QNDM quasi-probability density function $\mathcal{P}_{\rm QNDM}(\Delta)$ comprises only real numbers. Second, $\mathcal{P}_{\rm QNDM}(\Delta)$ is normalized to $1$, and this normalization comes uniquely from the classical contribution $\mathcal{P}_{cl}(\Delta)$. The first and second properties imply that the terms entering the quantum contribution $\mathcal{P}_{q}(\Delta)$ of $\mathcal{P}_{\rm QNDM}(\Delta)$ cancel out each other, such that some of them must be negative. The last property, which directly links with the violation of the macrorealism, is that the negativity of $\mathcal{P}_{\rm QNDM}(\Delta)$ appears whenever the system is in a superposition of the measurement observable’s eigenstates.
@@
Let us argue about the meaning of the necessary and sufficient conditions for the violation of MR. From the one side, since the negativity of $\mathcal{P}_{\rm QNDM}(\Delta)$ derives from contributions in $\mathcal{P}_{q}(\Delta)$, a necessary condition for the negativity of $\mathcal{P}_{\rm QNDM}(\Delta)$ is that the system is in a coherent superposition of the observable $\hat{A}$’s eigenstates at times $t_{0}$ or $t_{1}$ or $t_{2}$. On the other hand, the fact that the system is in a coherent superposition state is also a sufficient condition for the negativity of $\mathcal{P}_{\rm QNDM}(\Delta)$. This is because the negativity of $\mathcal{P}_{\rm QNDM}(\Delta)$ can derive only from $\mathcal{P}_{q}(\Delta)$, whose integral over all values of $\Delta$ is equal to zero: $\int d\Delta\mathcal{P}_{q}(\Delta)=0$. This constraint imposes that, whenever $\mathcal{P}_{q}(\Delta)\neq 0$, $\mathcal{P}_{\rm QNDM}(\Delta)$ must contain negative real terms.
@@
Two observations are now in order. First, by construction, $\mathcal{P}_{q}(\Delta)$ comprises real numbers that are the real parts of $P_{\rm QNDM}(k,j,m,i,l)$, complex numbers. This means that the sum of $P_{\rm QNDM}(k,j,m,i,l)$ could vanish even if the system is a coherent superposition of $\hat{A}$’s eigenstates, when the quasiprobabilities $P_{\rm QNDM}(k,j,m,i,l)$ are purely imaginary. To avoid this, it is sufficient to run the measurement protocol two more times, by performing the additional unitary transformation $\hat{U}_{\chi}\mathrel{\mathop{\ordinarycolon}}=\exp\{i\chi\hat{A}\}$ on the system — with $\chi$ a given angle — first at time $t_{0}$ and then at time $t_{1}$ in the second run. The transformation $\hat{U}_{\chi}$ just adds a phase in the system state and does not induce transitions between the eigenstates of the observable $\hat{A}$. Therefore, it does not change the fact that the system is in a superposition state and thus that it may satisfy the assumption of MRps. A proper choice of $\chi$ (even sampled by a probability distribution could be a good option) ensures that every pure imaginary amplitude encoded in $P_{\rm QNDM}(k,j,m,i,l)$ acquires a real part, resulting in a $\mathcal{P}_{q}(\Delta)\neq 0$.
@@
In addition, for practical implementations, the whole procedure for quantumness certification becomes more robust if the second and third measurements at times $t_{1}$ and $t_{2}$ are made over the measurement observable with slightly shifted coupling parameters: $\lambda\rightarrow\lambda^{\prime}=\lambda+\delta$ and $\lambda\rightarrow\lambda^{\prime\prime}=\lambda+\delta^{\prime}$. These measurements are realized by applying the unitary transformation $\hat{u}_{\delta}\mathrel{\mathop{\ordinarycolon}}=e^{i\frac{\lambda^{\prime}}{2}\hat{A}\otimes\hat{p}}$ and $\hat{u}_{\delta^{\prime}}\mathrel{\mathop{\ordinarycolon}}=e^{i\frac{\lambda^{\prime\prime}}{2}\hat{A}\otimes\hat{p}}$ to the system-detector. The advantages of applying $\hat{u}_{\delta},\hat{u}_{\delta^{\prime}}$ are to split the support of $\mathcal{P}_{\rm QNDM}(\Delta)$ in $\Delta$ and to remove the degeneracy among the different paths that are associated with the same value of $\Delta_{k,j,m,i,l}$. For example, for $i=j$, $l=m$ and $i\neq l$, the system is in a coherent superposition of states and $\Delta=a_{k}$. However, the same value of $\Delta$ is also associated to the classical path with $i=l$ and $j=m$. This fact could happen for binary observables (see also below) when ($a_{i}=a_{j}=1,\,a_{m}=a_{l}=-1$) or ($a_{i}=a_{j}=-1,\,a_{m}=a_{l}=1$). But if we split the support of $\mathcal{P}_{\rm QNDM}(\Delta)$, we can guarantee that both classical and quantum contributions are present in $\Delta$.
@@
Overall, the procedure for operationally certifying a violation of MR is given by the following three runs of experiments:
@@
If none of the resulting QNDM quasi-probability density functions $\mathcal{P}_{\rm QNDM}(\Delta)$ show negative regions, then MR is fulfilled, meaning that the system under scrutiny can be considered as classical according to the measurement observable $\hat{A}$. This means that the system has definite properties at all times, and it satisfies the MR conditions.
IV.3.1 Comparison with previous approaches
In the literature, there are already some papers that address the existence of a necessary and sufficient condition for the violation of MR. Among them, it is worth mentioning the ones by Clemente and Kofler [27, 28] and by Halliwell [29, 30]. In particular, in Ref. [30], the Author makes a quite accurate review of theoretical approaches that aim to identify MR violations.
In Ref. [30], Halliwell distinguishes between two approaches for identifying a violation of MR, denoted respectively as MR-weak and MR-strong, which are derived from two versions of NIM: the “weak” or “piecewise” NIM (NIM-pw) and the “strong” or “sequential” one (NIM-seq). NIM-pw is inherently satisfied in a Leggett–Garg measurement scheme, while NIM-seq is operationally replaced by the NSIT conditions, which are statistically equivalent.
As discussed in Ref. [30], the MR-weak approach is not always able to detect quantum coherences (or interferences) that are present in the initial state of the system or generated dynamically. By contrast, the MR-strong approach provides a much stronger framework for identifying MR violations, although it is not guaranteed that NIM-seq is satisfied for any choice of parameters entering the initial state and the measurement observable. In this regard, from [30] we quote the sentence “it is, therefore, necessary to adjust the initial state and measurement times (and perhaps other parameters too) to ensure that the NSIT conditions are satisfied”, whose analogue with the same meaning can be found in the example and summary table of Ref. [27]. This is because, for any given combination of initial state, system dynamics, and measurements, ensuring that the NSIT conditions are valid at all times (three times in our case study) requires solving a system of many parametric equations [27]. The size of this system grows with the dimension of the system under consideration. Moreover, it is not guaranteed that the system of equations always admits solutions that confirm or exclude violations of MR [30].
For quantum systems with many degrees of freedom or highly correlated dynamics, it may be difficult or even impossible to verify the validity of the NSIT conditions. Indeed, it is unknown whether there exists a set of parameters (initial state, evolution, and measurements) that leads to the violation of at least one NSIT condition in the presence of quantum coherence and/or correlations in the system state [27, 30]. This fact is clearly stated in Ref. [27], where it is written: “While our results suggest that an experimental demonstration of nonclassicalities requires either very precise measurements or a complex time evolution, a general proof of this trade-off (in terms of experimental control parameters) is still missing.”
Given this context, the QNDM approach is instead a viable, operational protocol to certify MR. Such an approach falls within the category of NIM-seq, as it satisfies all NSIT conditions by construction. Moreover, the QNDM protocol is designed to identify a violation of MRps, independently of the initial state, measurement observables, or time evolution of the system. These violations are certified by examining a single quantity: the negativity of the QNDM quasi-probability density function. In addition, the QNDM approach has already found successful application in many experimental scenarios [81, 82, 22, 83, 23, 84], as it only requires coupling the measured system with a detector via a phase operator that has broad experimental applicability. In particular, in Refs. [22, 23], QNDM have been employed to monitor the loss of coherence due to environmental effects in commercial IBM quantum processors. These tests have shown that a QNDM protocol can be realized even in a setting hardly affected by noise.
Similar ideas of using an additional quantum system to reconstruct features of a quantum system has been proposed also in the field of full-counting statistics [85, 86, 87], where the statistics of electrons flowing through a wire is measured. While it was known that the approach of full-counting statistics can lead to a quasi-probability distribution [87, 88, 55, 89, 90], implications in terms of macrorealism violation and the connection of the latter to the negativity of the quasi-probability distribution have not been fully understood. However, it is worth mentioning that the results derived using full-counting statistics led to several implementations in condensed matter and electronic circuits [91, 92, 93], thus bringing advancement in such fields.

### v2

@@ IV.1 Three-measurement scheme
The simplest and most used version of the Leggett-Garg inequalities (LGIs) [5] is based on a three measurement scheme [83, 6]. To discuss the relations between the QNDM protocol and LGIs, we extend the protocol discussed in Sec. III to the three-measurement scheme depicted in Fig. 1. For technical details, the reader can refer to Ref. [26].
@@
Let us thus consider a system that evolves under a unitary transformation in the time interval $0\leq t\leq\mathcal{T}$, and the system observable $\hat{A}$ is measured sequentially at times $t_{0}\leq t_{1}\leq t_{2}=\mathcal{T}$. As above, the measurements of $\hat{A}$ are performed by coupling the system with a detector that interacts through the operator $\hat{u}$. The system’s time evolution in between the measurements is governed by the unitary operators $\hat{U}_{k}\equiv\hat{U}(t_{k},t_{k-1})\otimes\mathbbm{1}$ with $k=1,2$. Hence, overall, the time evolution operator $\hat{U}_{tot}$ for the system-detector within the interval $t_{0},t_{2}]$ read as
@@
$P_{\rm QNDM}(k,j,m,i,l)=U_{2,kj}U_{1,ji}\,\rho^{0}_{il}U^{*}_{1,lm}U^{*}_{2,mk}={\rm Tr}\left[\hat{\Pi}_{k}\hat{U}_{2}\hat{\Pi}_{j}\hat{U}_{1}\hat{\Pi}_{i}\,\hat{\rho}^{0}\hat{\Pi}_{l}\hat{U}^{\dagger}_{1}\hat{\Pi}_{m}\hat{U}^{\dagger}_{2}\right]$ (23)
are the interference contributions for the system to arrive in the state $\mathinner{|{k}\rangle}$ following the superposition of paths $i\rightarrow j\rightarrow k$ and $l\rightarrow m\rightarrow k$ (see Fig. 1).
Again, we can separate the classical $\mathcal{P}_{cl}(\Delta)$ from the quantum $\mathcal{P}_{q}(\Delta)$ contributions of $\mathcal{P}_{\rm QNDM}(\Delta)$. The first ones are obtained when $m=j$ and $l=i$, i.e.,
$\mathcal{P}_{cl}(\Delta)=\sum_{i,j,k}P_{\rm QNDM}(k,j,j,i,i)\,\delta\left[\Delta-\Delta_{k,j,j,i,i}\right]$ (24)
that comprise positive probabilities. This is because $P_{\rm QNDM}(k,j,j,i,i)=P^{(0)}_{i}P^{(1)}_{i\rightarrow j}P^{(2)}_{j\rightarrow k}$, where $P^{(0)}_{i}=\rho^{0}_{ii}$, and $P^{(1)}_{i\rightarrow j}=|U_{1,ji}|^{2},P^{(2)}_{j\rightarrow k}=|U_{2,kj}|^{2}$ are the probabilities the system is initially in the state $\mathinner{|{i}\rangle}$ and then makes a transition from state $\mathinner{|{i}\rangle}$ to $\mathinner{|{j}\rangle}$ and from $\mathinner{|{j}\rangle}$ to $\mathinner{|{k}\rangle}$, respectively.
The quantum contributions $\mathcal{P}_{q}(\Delta)$ of $\mathcal{P}_{\rm QNDM}(\Delta)$, instead, are present when $m\neq j$ or $l\neq i$. Thus, they are associated with the coherent superposition of paths or trajectories in the space identified by the sequence of measurement outcomes [19, 26]. They formally read as
@@
IV.1.1 QNDM distribution implies unperturbed marginals
The quasi-probability density function in (22) has a significant mathematical property: the marginals of $\mathcal{P}_{\rm QNDM}(\Delta)$, over all the measurement outcomes but the ones at a single time, return the ‘unperturbed’ probability to measure the observable $\hat{A}$ given by the Born’s rule, without any intermediate measurement of the sequence affects the result.
@@ IV.1.1 QNDM distribution implies unperturbed marginals
$P(a_{k})=\sum_{i,l,j,m}P_{ND}(k,j,m,i,l)=\sum_{i,l,j,m}{\rm Tr}\left[\hat{\Pi}_{k}\hat{U}_{2}\hat{\Pi}_{j}\hat{U}_{1}\hat{\Pi}_{i}\,\hat{\rho}^{0}\hat{\Pi}_{l}\hat{U}^{\dagger}_{1}\hat{\Pi}_{m}\hat{U}^{\dagger}_{2}\right]={\rm Tr}\left[\hat{\Pi}_{k}\hat{U}\hat{\rho}^{0}\hat{U}^{\dagger}\right]={\rm Tr}\left[\hat{\Pi}_{k}\,\hat{\rho}(t_{2})\right],$ (26)
where $\hat{U}=\hat{U}_{2}\hat{U}_{1}$ and $\hat{\rho}(t_{2})=\hat{U}\hat{\rho}^{0}\hat{U}^{\dagger}$. Therefore, $P(a_{k})$ is the probability to record the measurement outcome $a_{k}$ at time $t_{2}$ as given by Born’s rule [6]. We also compute the single-time probabilities $P(a_{i})$, $P(a_{l})$ for the initial measurement at time $t_{0}=0$:
@@
$=$ $\frac{1}{2}\sum_{l,j,m,k}\Big\{{\rm Tr}\left[\hat{\Pi}_{k}\hat{U}_{2}\hat{\Pi}_{j}\hat{U}_{1}\hat{\Pi}_{i}\,\hat{\rho}^{0}\hat{\Pi}_{l}\hat{U}^{\dagger}_{1}\hat{\Pi}_{m}\hat{U}^{\dagger}_{2}\right]+{\rm Tr}\left[\hat{\Pi}_{k}\hat{U}_{2}\hat{\Pi}_{j}\hat{U}_{1}\hat{\Pi}_{l}\,\hat{\rho}^{0}\hat{\Pi}_{i}\hat{U}^{\dagger}_{1}\hat{\Pi}_{m}\hat{U}^{\dagger}_{2}\right]\Big\}={\rm Tr}\left[\hat{\Pi}_{i}\,\hat{\rho}^{0}\right]$
@@
both these probabilities satisfy Born’s rule, as well as $P(a_{k})$. Moreover, for the measurement at time $t_{1}$, there are two possible realizations over the indices $j$ and $m$, so that
$P(a_{j})=\frac{1}{2}\sum_{i,l,m,k}\Big\{P_{ND}(k,j,m,i,l)+P_{ND}(k,m,j,i,l)\Big\}={\rm Tr}\left[\hat{\Pi}_{j}\hat{U}_{1}\hat{\rho}^{0}\hat{U}^{\dagger}_{1}\right]={\rm Tr}\left[\hat{\Pi}_{j}\,\hat{\rho}(t_{1})\right],$ (29)
@@
which is the probability of recording the outcome $a_{j}$ after the system has evolved to the state $\hat{\rho}(t_{1})=\hat{U}_{1}\,\hat{\rho}^{0}\hat{U}^{\dagger}_{1}$, without the initial measurement at time $t=0$ has affected the system’s state.
@@
Finally, summing the QNDM quasi-probabilities $P_{\rm QNDM}(k,j,m,i,l)$ over one of the possible three-index sequences returns the Kirkwood-Dirac quasi-probabilities at two-times [54, 25, 68], apart from the 3-tuples $(j,m,k)$ and $(i,l,k)$ that give rise to a single-time probability. By definition, also Kirkwood-Dirac quasi-probabilities return marginal distributions that are not perturbed by any intermediate measurement performed previously.
IV.2 Theoretical basis for quantumness and classicality certification
With a few exceptions of negligible statistical significance (corresponding to isolated points in a large parameter space) that can be cured as discussed in the next section IV.3, we can prove that $\mathcal{P}_{q}(\Delta)\neq 0$ always identifies whether the system’s state is in a superposition of the $\hat{A}$’s eigenstates.
@@ IV.2 Theoretical basis for quantumness and classicality certification
The quantumness certification stems from the following statement: the presence of negative regions in the QNDM quasi-probability density function $\mathcal{P}_{\rm QNDM}(\Delta)$ is a necessary and sufficient condition for the violation of the MRps [26]. In other words, a system is in a coherent superposition of the measured observable’s eigenstates if and only if $\mathcal{P}_{\rm QNDM}(\Delta)$ exhibits negativity, i.e., it has negative contributions. The fact that the measured system is in a superposition state violates the MRps and, consequently, the MR.
@@
This result is based on the following three main properties of the QNDM protocol, derived in Ref. [26]. First, the QNDM quasi-probability density function $\mathcal{P}_{\rm QNDM}(\Delta)$ comprises only real numbers. Second, $\mathcal{P}_{\rm QNDM}(\Delta)$ is normalized to $1$, and this normalization comes uniquely from the classical contribution $\mathcal{P}_{cl}(\Delta)$. The first and second properties imply that the terms entering the quantum contribution $\mathcal{P}_{q}(\Delta)$ of $\mathcal{P}_{\rm QNDM}(\Delta)$ cancel out each other, such that some of them must be negative.
@@
The negativity of $\mathcal{P}_{\rm QNDM}(\Delta)$ appears whenever the system is in a superposition of the measurement observable $\hat{A}$’s eigenstates, in at least one measurement time among $t_{0},t_{1},t_{2}$. This is because the negativity of $\mathcal{P}_{\rm QNDM}(\Delta)$ univocally means that $\mathcal{P}_{q}(\Delta)\neq 0$, which occurs if and only if the system’s state is in a superposition of the $\hat{A}$’s eigenstates. Hence, since the integral of $\mathcal{P}_{q}(\Delta)$ over all values of $\Delta$ is equal to zero ($\int d\Delta\mathcal{P}_{q}(\Delta)=0$), $\mathcal{P}_{\rm QNDM}(\Delta)$ must contain negative real terms whenever $\mathcal{P}_{q}(\Delta)\neq 0$.
The positivity of the QNDM quasi-probability density function gives us complementary information about the system’s behaviours, and can thus be used as a witness for classicality certification. In fact, since the positivity of the QNDM distribution $\mathcal{P}_{\rm QNDM}(\Delta)$ implies the absence of quantum superposition, it entails that the system’s states at the measurement times $t_{0},t_{1},t_{2}$ are diagonal along the basis that spectrally decomposes the observable $\hat{A}$ at such times [25]. Hence, $\mathcal{P}_{q}(\Delta)=0$ and $\mathcal{P}_{\rm QNDM}(\Delta)=\mathcal{P}_{\rm cl}(\Delta)$. As discussed in the previous section, $\mathcal{P}_{\rm cl}(\Delta)$ is obtained from implementing the TPM scheme [25], which is realized by a sequence of projective measurements of $\hat{A}$ at times $t_{0},t_{1},t_{2}$. In this sense, $\mathcal{P}_{\rm cl}(\Delta)$ is denoted as being classical. In the conditions where the system’s states are diagonal from the positivity of the QNDM distribution, the TPM protocol does not entail any perturbation of the outcomes’ statistics from measuring $\hat{A}$. This means that the TPM distribution suffices to describe the statistics of measurement outcomes at times $t_{0},t_{1},t_{2}$, as it is able to return the unperturbed value of the corresponding statistical moments [25]. The unperturbed moments are obtained because the system’s states are diagonal along the observable’s basis at the measurement times. Therefore, no quantum coherence and/or correlation is erased on such a basis, meaning that a classical path must be followed (red dashed curve in Fig. 1).
Alternatively, the concepts about classicality certification discussed in the latter paragraph can be figured out by noting that each term $P_{\rm QNDM}(k,j,j,i,i)$ comprising the positive distribution $\mathcal{P}_{\rm cl}(\Delta)=\sum_{i,j,k}P_{\rm QNDM}(k,j,j,i,i)\,\delta[\Delta-\Delta_{k,j,j,i,i}]$ in (24) is associated with the averages and correlators $\langle\hat{A}_{i}\rangle=\sum_{a_{i}}a_{i}\,P_{i}^{(0)}(a_{i})$, $\langle\hat{A}_{j}\rangle=\sum_{a_{i},a_{j}}a_{j}\,P_{i}^{(0)}(a_{i})\,P_{i\to j}^{(1)}(a_{j}|a_{i})$, $\langle\hat{A}_{i}\hat{A}_{j}\rangle=\sum_{a_{i},a_{j}}a_{i}a_{j}\,P_{i}^{(0)}(a_{i})\,P_{i\to j}^{(1)}(a_{j}|a_{i})$, $\langle\hat{A}_{i}\hat{A}_{j}\hat{A}_{k}\rangle=\sum_{a_{i},a_{j},a_{k}}a_{i}a_{j}a_{k}\,P_{i}^{(0)}(a_{i})\,P_{i\to j}^{(1)}(a_{j}|a_{i})\,P_{j\to k}^{(2)}(a_{k}|a_{j})$ and so on, built over the classical path $i\rightarrow j\rightarrow k$. In particular,
$P_{\rm QNDM}(k,j,j,i,i)=\frac{1}{8}\Big[1+a_{i}\langle\hat{A}_{i}\rangle+a_{j}\langle\hat{A}_{j}\rangle+a_{k}\langle\hat{A}_{k}\rangle+a_{i}a_{j}\langle\hat{A}_{i}\hat{A}_{j}\rangle+a_{i}a_{k}\langle\hat{A}_{i}\hat{A}_{k}\rangle++a_{j}a_{k}\langle\hat{A}_{j}\hat{A}_{k}\rangle+a_{i}a_{j}a_{k}\langle\hat{A}_{i}\hat{A}_{j}\hat{A}_{k}\rangle\Big]$ (30)
that is in the same form as a generic three-time probabilities as discussed in Ref. [84]. Therefore, we can conclude that the positivity of $P_{\rm QNDM}(k,j,j,i,i)$ can be used to identify the classical behaviour of the quantum system.
IV.3 Operational procedure for quantumness certification
At this point, we can show how the QNDM quasi-probability density function can be extracted with a feasible and operational procedure that always guarantees the certification of quantum features in a system’s state. To do this, we explicitly take into account the few exceptions whereby $\mathcal{P}_{q}(\Delta)$ is not able to identify superpositions of the $\hat{A}$’s eigenstates in the system’s state.
@@ IV.3 Operational procedure for quantumness certification
First, by construction, $\mathcal{P}_{q}(\Delta)$ comprises real numbers that are the real parts of $P_{\rm QNDM}(k,j,m,i,l)$, complex numbers. This means that the sum of $P_{\rm QNDM}(k,j,m,i,l)$ could vanish even if the system is a coherent superposition of $\hat{A}$’s eigenstates, when the quasi-probabilities $P_{\rm QNDM}(k,j,m,i,l)$ are purely imaginary. To avoid this, it is sufficient to run the measurement protocol two more times, by performing the additional unitary transformation $\hat{U}_{\chi}\mathrel{\mathop{\ordinarycolon}}=\exp\{i\chi\hat{A}\}$ on the system, first at time $t_{0}$ and then at time $t_{1}$ in the second run. The transformation $\hat{U}_{\chi}$ just adds a phase in the system’s state and does not induce transitions between the eigenstates of the observable $\hat{A}$. Therefore, it does not change the fact that the system is in a superposition state. A proper choice of $\chi$ (even sampled by a probability distribution could be a good option) ensures that every pure imaginary contribution encoded in $P_{\rm QNDM}(k,j,m,i,l)$ acquires a real part, which results in $\mathcal{P}_{q}(\Delta)\neq 0$.
@@
In addition, the whole procedure for quantumness certification becomes more robust if the second and third measurements at times $t_{1}$ and $t_{2}$ are made over the measurement observable with slightly shifted coupling parameters: $\lambda\rightarrow\lambda^{\prime}=\lambda+\delta$ and $\lambda\rightarrow\lambda^{\prime\prime}=\lambda+\delta^{\prime}$. These measurements are realized by applying the unitary transformation $\hat{u}_{\delta}\mathrel{\mathop{\ordinarycolon}}=e^{i\frac{\lambda^{\prime}}{2}\hat{A}\otimes\hat{p}}$ and $\hat{u}_{\delta^{\prime}}\mathrel{\mathop{\ordinarycolon}}=e^{i\frac{\lambda^{\prime\prime}}{2}\hat{A}\otimes\hat{p}}$ to the system-detector. The advantages of using $\hat{u}_{\delta},\hat{u}_{\delta^{\prime}}$ are to split the support of $\mathcal{P}_{\rm QNDM}(\Delta)$ in $\Delta$, and to remove the degeneracy among the different paths associated with the same value of $\Delta_{k,j,m,i,l}$. For example, for $i=j$, $l=m$, and $i\neq l$, the system is in a coherent superposition of states and $\Delta=a_{k}$. However, the same value of $\Delta$ is also associated with the classical path with $i=l$ and $j=m$. This fact could happen for binary observables when ($a_{i}=a_{j}=1,\,a_{m}=a_{l}=-1$) or ($a_{i}=a_{j}=-1,\,a_{m}=a_{l}=1$). But if we split the support of $\mathcal{P}_{\rm QNDM}(\Delta)$, we can guarantee that both classical and quantum contributions are present in $\Delta$.
@@
Overall, the procedure for certifying a violation of MRps (and thus of MR) is given by the following three runs of experiments:
@@
If none of the resulting QNDM quasi-probability density functions $\mathcal{P}_{\rm QNDM}(\Delta)$ show negative regions, then MRps is fulfilled.